OpenStudy (anonymous):

Nathan stores coffee powder in a steel container with a capacity of 68.75 cubic inches. He uses a glass container to store sugar. The two containers are congruent cylinders. If the glass container has a height of 3.5 inches, what is its radius? Answer 2.5 inches 3.5 inches 5.0 inches 7.0 inches

OpenStudy (anonymous):

\[pi r^2 h =68.72\] r^2=68.72/ pi h if u substtut the gvn valus r=2.5

OpenStudy (anonymous):

thanks